After measuring the concentration of chlorophyll, you find that the concentration in your chloroplast isolate is 4.0 µg/µl. Find the volume (in µl) of chloroplast suspension need to have 20 µg of chlorophyll in your final reaction. Your final reaction has a volume of 1 ml.
Answer #1

Given- 4 ug/ul

Final volume =?

Final amount =20 ug

Chlorophyll amount has been increased by a factor of - 20/4 = 5

Volume will also be increased by same factor

So volume of stock required = 1 × 5 = 5 ul

So if 1 ul contains 4 ug then, 5 ul would have 20 ug.

So take 5 ul of solution and dissolve in final reaction volume that us 1 ml.